# Interpreting Language Reward Models via Contrastive Explanations

### v1

@@ 1 Introduction
A complementary approach for achieving better understandings of complex machine learning models is to apply explainable AI (XAI) techniques, which aim to uncover reasons behind their predictions. In addition to the complexity of the language domain, a particular challenge for understanding RMs is the lack of direct meaning in their predicted rewards. RMs are typically trained and evaluated on binary response comparisons (Ouyang et al., 2022) using the preference model of Bradley & Terry (1952), which converts the RM’s scalar reward for each response into a probabilistic preference for one response or the other. Outside of such a comparison, the meaning of the output scalars themselves is unclear, but traditional XAI techniques usually require being able to interpret the output values as part of the explanation. Recent studies propose to provide more information in RM outputs by decomposing one reward into multiple evaluation aspects like correctness, relevance and verbosity (Wang et al., 2023; Go et al., 2024; Wang et al., 2024a). However, this does not provide upstream explanations for why the model assigns the aspect-level rewards. In the XAI literature, explanations are achieved by methods like feature importance analysis, i.e. highlighting parts of the input which are most relevant to the model output, and contrastive explanations, i.e. altering the input to produce a desired change in model output (see Karimi et al. (2023); Aryal & Keane (2023) for recent overviews). To our knowledge, no such XAI techniques have been explored for RMs.
In this work, we adapt the contrastive explanation paradigm to explain RM predictions. The two most common variants are counterfactual (CF) explanations (Wachter et al., 2017) and semifactual (SF) explanations (Kenny & Keane, 2021), which are defined for classification settings as follows. A CF is a minimally perturbed input that yields an alternative prediction from the model, providing information about a what if scenario: had the input been different, the prediction would have been different. An SF is a maximally perturbed input for which the prediction does not change, indicating how even if the input were different, the prediction result would be the same. Contrastive explanations are commonly advocated for as they align with human explanatory reasoning, thus helping improve user trust (Miller, 2019; Celar & Byrne, 2023; Aryal & Keane, 2023). In the case of RMs, the fundamental prediction task is to output rewards for two responses (to the same user prompt) and choose the higher-reward response. Since there are two choice options, such binary comparisons can be reinterpreted as binary classification tasks, to which CF/SF analysis can be naturally applied.
@@
We obtain a global summary of an RM’s sensitivity to each evaluation attribute by aggregating CFs and SFs over multiple comparisons, as similarly done in, e.g., global recourse (Rawal & Lakkaraju, 2020) and global feature importance (Ribeiro et al., 2016). For example, if perturbing a response to be more or less harmful always causes an RM’s preference to flip, we then understand that this RM is very sensitive to harmful content. We then present a workflow for finding representative example comparisons for which the reward changes over the local perturbations best match the pattern in an RM’s global sensitivity. Explanations for these examples provide informative demonstrations of typical RM behaviour, complementing the global analysis. The contributions of our work include:
@@ 2.3 Generating Perturbed Responses
Desiderata. For explanations to provide meaningful insight, the perturbed responses should be well-formed sentences which are grammatically correct and semantically interpretable. This makes naïve methods based on randomly replacing characters, tokens or words unsuitable. Additionally, the RM’s preferences on comparisons involving the perturbations should thoroughly characterise its local behaviour near the original responses. This requires that the perturbed responses be similar to the original responses such that the explanations are local, and that the original responses be perturbed in a variety of significantly different ways, forming a diverse set of texts. Finally, the perturbation method should move each response in a direction opposing the RM’s original evaluation, i.e. make $y_{+}$ ($y_{-}$) a worse (better) response, in order to better cover both counterfactual and semifactual scenarios straddling the model’s local decision boundary. These highlighted desiderata are consistent with those previously suggested for CFs for text classification (Nguyen et al., 2024).
@@
Response perturbation via LLMs prompting. Generative language models have proven effective at producing well-formed sentences for use in explanation (Sen et al., 2023) and RM evaluation (Zeng et al., 2024; Park et al., 2024), and we follow this precedent here. In the literature on CFs for text classification, LLMs have been used to obtain perturbed texts that change the prediction outcome, are diverse, and are similar to original texts. Earlier works by Yang et al. (2020) and Wu et al. (2021) rely on masked language models and GPT-2 (Radford et al., 2019) to perform elementary perturbations like word insertion, deletion and replacement. More recently, various prompting strategies have been proposed to utilise publicly available LLMs like GPT-4 (Achiam et al., 2023) and Llama (AI@Meta, 2024) for perturbing texts to obtain CFs (Gat et al., 2024; Cheng et al., 2024; Bhattacharjee et al., 2024). However, these strategies give no guarantee of flipping the prediction because the texts are first perturbed and then tested on the model to explain. Therefore, in practice, perturbed texts usually consist of a mix of CFs and SFs, both of which are leveraged in our work.
@@
Attribute-conditioned prompting strategy. To satisfy the desiderata of locality and diversity, we propose a two-step LLM prompting strategy to generate ${\mathbb{Y}}_{+}$ and ${\mathbb{Y}}_{-}$. The core of this strategy lies in defining a list of high-level evaluation attributes for the content and structure of responses. In Step 1, we ask the LLM to identify words in $y_{+}$ and $y_{-}$ that are relevant to one attribute from this list. Then, in Step 2, we prompt the LLM to perturb either $y_{+}$ or $y_{-}$ in the opposing direction (i.e. making $y_{+}$ worse; making $y_{-}$ better) along that attribute. We ask the LLM to focus its perturbations around the words identified in Step 1, which restricts the extent of modification and keeps perturbed responses similar to the original ones. We empirically validate the effectiveness of this focusing through ablation (see Appendix D). In our experiments, ${\mathbb{Y}}_{+}$ and ${\mathbb{Y}}_{-}$ both contain 15 perturbed responses, each associated with one attribute from the following list: avoid-to-answer, appropriateness, assertiveness, clarity, coherence, complexity, correctness, engagement, harmlessness, helpfulness, informativeness, neutrality, relevance, sensitivity, verbosity. See Appendices A and B for details about attributes and prompts.
@@ 3 Quantitative Evaluation
Datasets. We use HelpSteer2 (hs2) (Wang et al., 2024b), HH-RLHF-helpful, and HH-RLHF-harmless [^1] HH-RLHF-helpful and HH-RLHF-harmless are two branches of the HH-RLHF dataset. We treat them separately because they have distinct characteristics. (Bai et al., 2022). We focus on explaining single-turn conversations in our experiments. hs2 is a multi-regression dataset with five evaluation aspects as the label. We filter hs2 for pairs of responses where the preference is clear, i.e. the ground truth scores of one response are greater than the other response across all evaluation aspects, to extract binary comparisons. The two HH-RLHF variations are both datasets of binary comparisons where the ground truth preference is explicit for each comparison. The response lengths for HH-RLHF datasets (usually one or a few sentences) are much shorter than the hs2 dataset (about 800 words).
@@
Reward Models. We generate explanations for three open source RMs from the OpenAssistant community. DeBERTa-large (v1) [^2] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large and DeBERTa-large-v2333https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2, comprising 0.4 billion parameters, are RMs fine-tuned from the DeBERTa model (He et al., 2021) using different datasets (see their model cards for more details). pythia-1.4b-epoch-2.5444https://huggingface.co/OpenAssistant/oasst-rm-2.1-pythia-1.4b-epoch-2.5 is a larger model with 1.4 billion parameters.
@@
Metrics. We use a set of metrics derived from the desiderata in Section 2.3. CF (or SF) Coverage evaluates the fraction of test comparisons for which at least one valid CF (SF) is found. Having both high CF and SF coverage indicates that an explainer is better at obtaining a balanced mix of CFs and SFs. To measure the similarity between perturbed responses and their associated original response, we use two distance metrics between sentences. Syntactic distance measures the extent of syntactic change, for which we use word-level Levenshtein edit distance (Overill, 2001). Semantic distance, a common metric for measuring differences in meaning, is the cosine distance between the Sentence-BERT encodings of the two text pieces (Reimers & Gurevych, 2019). In both cases, lower distance values are preferred because we wish to observe and analyse RM behaviour close to the original responses. Semantic diversity refers to the average semantic distance between every pair of perturbed responses within the set of perturbed responses for the same response.
@@
Baselines. We compare our method to two baselines for generating perturbed responses, from which CFs and SFs are identified using the same categorisation procedure (Section 2.2). We adapt Polyjuice (PJ) by Wu et al. (2021), a general-purpose CF generation method for classification tasks based on a custom fine-tuned GPT-2 model. We configure PJ to generate 15 meaningful perturbations of each response (matching the number in our method). Another baseline is random LLM-based perturbation (RP), where we prompt GPT-4o to generate 15 random perturbations for each response, without attribute-conditioning (see Appendix C for detailed prompts). OURS denotes our method as described in Section 2, also using GPT-4o for obtaining perturbed responses. Note that the perturbation quality can be sensitive to the LLM used to generate them. We choose the best-performing one here and refer to Bhattacharjee et al. (2024); Sen et al. (2023) for further analysis.

### v2

@@ 1 Introduction
A complementary approach for achieving better understandings of complex machine learning models is to apply explainable AI (XAI) techniques, which aim to uncover reasons behind their predictions. In addition to the complexity of the language domain, a particular challenge for understanding RMs is the lack of direct meaning in their predicted rewards. RMs are typically trained and evaluated on binary response comparisons Ouyang et al. (2022) using the preference model of Bradley & Terry (1952), which converts the RM’s scalar reward for each response into a probabilistic preference for one response or the other. Outside of such a comparison, the meaning of the output scalars themselves is unclear, but traditional XAI techniques usually require being able to interpret the output values as part of the explanation. Recent studies propose to provide more information in RM outputs by decomposing one reward into multiple evaluation aspects like correctness, relevance and verbosity (Wang et al., 2023; Go et al., 2024; Wang et al., 2024a). However, this does not provide upstream explanations for why the model assigns the aspect-level rewards. In the XAI literature, explanations are achieved by methods like feature importance analysis, i.e. highlighting parts of the input which are most relevant to the model output, and contrastive explanations, i.e. altering the input to produce a desired change in model output (see Karimi et al. (2023); Aryal & Keane (2023) for recent overviews). To our knowledge, no such XAI techniques have been explored for RMs.
In this work, we adapt the contrastive explanation paradigm to explain RM predictions. The two most common variants are counterfactual (CF) explanations (Wachter et al., 2017) and semifactual (SF) explanations (Kenny & Keane, 2021), which are defined for classification settings as follows. A CF is a perturbed input that yields an alternative prediction from the model, providing information about a what if scenario: had the input been different, the prediction would have been different. An SF is a perturbed input for which the prediction does not change, indicating how even if the input were different, the prediction result would be the same. Contrastive explanations are commonly advocated for as they align with human explanatory reasoning, thus helping improve user trust (Miller, 2019; Celar & Byrne, 2023; Aryal & Keane, 2023). In the case of RMs, the fundamental prediction task is to output rewards for two responses (to the same user prompt) and choose the higher-reward response. Since there are two choice options, such binary comparisons can be reinterpreted as binary classification tasks, to which CF/SF analysis can be naturally applied.
@@
We obtain a global summary of an RM’s sensitivity to each evaluation attribute by aggregating CFs and SFs over multiple comparisons, as similarly done in other settings such as global recourse (Rawal & Lakkaraju, 2020) and global feature importance (Ribeiro et al., 2016). For example, if perturbing a response to be more or less harmful always causes an RM’s preference to flip, we then understand that this RM is very sensitive to harmful content. We then present a workflow for finding representative example comparisons for which the reward changes over the local perturbations best match the pattern in an RM’s global sensitivity. Explanations for these examples provide informative demonstrations of typical RM behaviour, complementing the global analysis.
We summarise our contributions as follows:
@@ 2.3 Generating Perturbed Responses
For explanations to provide meaningful insight, the perturbed responses should be well-formed sentences which are grammatically correct and semantically interpretable. This makes naïve methods based on randomly replacing characters, tokens or words unsuitable. Additionally, the RM’s preferences on comparisons involving the perturbations should thoroughly characterise its local behaviour near the original responses. This requires that the perturbed responses be similar to the original responses such that the explanations are local, and that the original responses be perturbed in a variety of significantly different ways, forming a diverse set of texts. Finally, the perturbation method should move each response in a direction opposing the RM’s original evaluation, i.e. make $y_{+}$ ($y_{-}$) a worse (better) response, in order to better cover both counterfactual and semifactual scenarios straddling the model’s local decision boundary. These highlighted desiderata are consistent with those previously suggested for CFs for text classification (Nguyen et al., 2024).
Response perturbation via LLMs prompting
@@
Generative language models have proven effective at producing well-formed sentences for use in explanation (Sen et al., 2023) and RM evaluation (Zeng et al., 2024; Park et al., 2024), and we follow this precedent here. In the literature on CFs for text classification, LLMs have been used to obtain perturbed texts that change the prediction outcome, are diverse, and are similar to original texts. Earlier works by Yang et al. (2020) and Wu et al. (2021) rely on masked language models and GPT-2 (Radford et al., 2019) to perform elementary perturbations like word insertion, deletion and replacement. More recently, various prompting strategies have been proposed to utilise publicly available LLMs like GPT-4 (Achiam et al., 2023) and Llama (AI@Meta, 2024) for perturbing texts to obtain CFs (Gat et al., 2024; Cheng et al., 2024; Bhattacharjee et al., 2024). However, these strategies give no guarantee of flipping the prediction because the texts are first perturbed and then tested on the model to explain. Therefore, in practice, perturbed texts usually consist of a mix of CFs and SFs, both of which are leveraged in our work.
Attribute-conditioned prompting strategy
@@
To satisfy the desiderata of locality and diversity, we propose a two-step LLM prompting strategy to generate ${\mathbb{Y}}_{+}$ and ${\mathbb{Y}}_{-}$. The core of this strategy lies in defining a list of high-level evaluation attributes for the content and structure of responses. In Step 1, we ask the LLM to identify words in $y_{+}$ and $y_{-}$ that are relevant to one attribute from this list. Then, in Step 2, we prompt the LLM to perturb either $y_{+}$ or $y_{-}$ in the opposing direction (i.e. making $y_{+}$ worse; making $y_{-}$ better) along that attribute. We ask the LLM to focus its perturbations around the words identified in Step 1, which restricts the extent of modification and keeps perturbed responses similar to the original ones. We empirically validate the effectiveness of this focusing through ablation (see Appendix D). In our experiments, ${\mathbb{Y}}_{+}$ and ${\mathbb{Y}}_{-}$ both contain 15 perturbed responses, each associated with one attribute from the following list: avoid-to-answer, appropriateness, assertiveness, clarity, coherence, complexity, correctness, engagement, harmlessness, helpfulness, informativeness, neutrality, relevance, sensitivity, verbosity. See Appendix A for descriptions and justifications of these attributes, and Appendix B for further details on prompts.
@@ 3 Quantitative Evaluation
We use HelpSteer2 (hs2) (Wang et al., 2024b), HH-RLHF-helpful, and HH-RLHF-harmless [^1] HH-RLHF-helpful and HH-RLHF-harmless are two branches of the HH-RLHF dataset. We treat them separately because they have distinct characteristics. (Bai et al., 2022). We focus on explaining single-turn conversations in our experiments. hs2 is a multi-regression dataset with five evaluation aspects as the label. We filter hs2 for pairs of responses where the preference is clear, i.e. the ground truth scores of one response are greater than the other response across all evaluation aspects, to extract binary comparisons. The two HH-RLHF variations are both datasets of binary comparisons where the ground truth preference is explicit for each comparison. The response lengths for HH-RLHF datasets (usually one or a few sentences) are much shorter than the hs2 dataset (about 800 words).
Reward Models
@@
We generate explanations for three open source RMs from the OpenAssistant community. DeBERTa-large (v1) [^2] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large and DeBERTa-large-v2333https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2, comprising 0.4 billion parameters, are RMs fine-tuned from the DeBERTa model (He et al., 2021) using different datasets (see their model cards for more details). pythia-1.4b-epoch-2.5444https://huggingface.co/OpenAssistant/oasst-rm-2.1-pythia-1.4b-epoch-2.5 is a larger model with 1.4 billion parameters.
Metrics
@@
We use a set of metrics derived from the desiderata in Section 2.3. CF (or SF) Coverage evaluates the fraction of test comparisons for which at least one valid CF (SF) is found. Having both high CF and SF coverage indicates that an explainer is better at obtaining a balanced mix of CFs and SFs. To measure the similarity between perturbed responses and their associated original response, we use two distance metrics between sentences. Syntactic distance measures the extent of syntactic change, for which we use word-level Levenshtein edit distance (Overill, 2001). Semantic distance, a common metric for measuring differences in meaning, is the cosine distance between the Sentence-BERT encodings of the two text pieces (Reimers & Gurevych, 2019). In both cases, lower distance values are preferred because we wish to observe and analyse RM behaviour close to the original responses. Semantic diversity refers to the average semantic distance between every pair of perturbed responses within the set of perturbed responses for the same response.
Baselines
@@
We compare our method to two baselines for generating perturbed responses, from which CFs and SFs are identified using the same categorisation procedure (Section 2.2). We adapt Polyjuice (PJ) by Wu et al. (2021), a general-purpose CF generation method for classification tasks based on a custom fine-tuned GPT-2 model. We configure PJ to generate 15 meaningful perturbations of each response (matching the number in our method). Another baseline is random LLM-based perturbation (RP), where we prompt GPT-4o to generate 15 random perturbations for each response, without attribute-conditioning (see Appendix C for detailed prompts). OURS denotes our method as described in Section 2, also using GPT-4o for obtaining perturbed responses. Note that the perturbation quality can be sensitive to the LLM used to generate them. We choose the best-performing one here and refer to Bhattacharjee et al. (2024); Sen et al. (2023) for further analysis.